Alpha-Tech, a rapidly growing distributor of electronic components, is formulating its plans for 20x5. Carol Jones, the firms marketing 28x5 Forecasted Sales s 7,580 8,509 7,580 18,000 11,008 12,580 February April 13,580 14,500 14,5e0 13,580 15,500 following information will be used in preparing the cash flow projection. erences s ex 40 percent two months after is Alpha-Techs largest expenditure, and each months cost of goods sold is estimated to month after the sale, and the . The be 35 percent of sales. Seventy percent of the parts are received by Alpha-Tech one month prior to sale, and 30 percent are received during the Historically 70 percent of accounts payable has been paid one month after receipt of the purchased components, and the of sale 30 percent has been paid two months after receipt Hourty wages and fringe benefits, estimated to be 30 percent of the current months sales, are paid in the month incurred . General and trative expenses are projected to be $15,720.000 for the year. The break of these expenses is presented All cash expenditures are paid uniformly throughout the year, except the property taxes, which are paid in four equal installments at the end of each quarter 3,6a0 2,500 1,600 Insurance utilities Depreciation Total $15,728
$3,100,000. The c . Alpha-Tech of is less than $520,000 at the end of each month, to Alpha-Tech has no short-term debt as of April 1, 20x5 . quarter of 20x5. For simplicity, ignore any interest expense associated with borrowing. (Negative amounts should be indicated by a minus sign) eferences Total receipts
Answer #1

Alpha Tech Cash Budget( In Thousands) For the Second Quarter of 20x5 Ma June Beginning Balance Collections: 520 5201,519 Febr

The opening balance for the month of April is $520,000, I have taken 520 as the figures are mentioned in 1000.

The sales for each month are mentioned.

60% of the billings are collected the month after the sale and 40% two months after the sale, thus the receipts for the month of April will be 60% of March sales 7500 + 40% of Feb sales 8500 which is 4500+3400 = 7900

Total cash available will be total receipts plus opening balance 520+7900 = 8420

Each months cost of goods sold is estimated to be 35% of sales, thus for april the COGS is 0.35*10,000 = 3500

70% of the Accounts payable are paid one month after the receipt of the purchased components and the remaining 30% two months after receipt. Thus 70% of March COGS 2625+ 30% of Feb COGS = 0.70*2625+0.30*2975 = 1837.5+892.5=2730

Wages 30% of the april sales are paid in april = 0.30*10,000 = 3000

All cash expenses are paid uniformly throughout the year except property taxes, thus the total expense excluding property tax is 15,720-1360= 14360/12= 1196.666

Income tax payments are made at the beginning of each calender quarter based on income of the prior quarter, thus 3100*0.40 = 1240

Thus the total disbursement will be the total of all from Accounts payable till Income tax

Cash balance will be Cash available - disbursements which is less than 520, thus the company will have to borrow 267

All the following months calculation is in the same way and only for the month of June there is a property tax expense.